### v1

# Coarse-Grained Boltzmann Generators

Abstract
Sampling equilibrium molecular configurations from the Boltzmann distribution is a longstanding challenge. Boltzmann Generators (BGs) address this by combining exact-likelihood generative models with importance sampling, but their practical scalability is limited. Meanwhile, coarse-grained surrogates enable the modeling of larger systems by reducing effective dimensionality, yet often lack the reweighting process required to ensure asymptotically correct statistics. In this work, we propose Coarse-Grained Boltzmann Generators (CG-BGs), a principled framework that unifies scalable reduced-order modeling with the exactness of importance sampling. CG-BGs act in a coarse-grained coordinate space, using a learned potential of mean force (PMF) to reweight samples generated by a flow-based model. Crucially, we show that this PMF can be efficiently learned from rapidly converged data via force matching. Our results demonstrate that CG-BGs faithfully capture complex interactions mediated by explicit solvent within highly reduced representations, establishing a scalable pathway for the unbiased sampling of larger molecular systems.
Machine Learning, ICML
1 Introduction
Accurately sampling molecular configurations from the Boltzmann distribution is a fundamental challenge in statistical physics (Chandler, 1987). Access to these samples is essential for computing thermodynamic observables, including free energies (Chipot and Pohorille, 2007). For molecular systems, the high dimensionality renders direct evaluation of the partition function intractable, forcing a reliance on simulation methods such as Molecular Dynamics (MD) or Markov Chain Monte Carlo (MCMC) (Frenkel and Smit, 2002). However, these approaches often struggle with the rugged, multi-modal structure of potential energy landscapes. High free-energy barriers trap trajectories in metastable states, leading to highly correlated samples and slow convergence (Lindorff-Larsen et al., 2011). Despite extensive progress in enhanced sampling methods (Zhu et al., 2025; Hénin et al., 2022) (e.g., umbrella sampling (Torrie and Valleau, 1977), metadynamics (Laio and Gervasio, 2008)) and coarse-graining (Noid, 2013; Saunders and Voth, 2013), efficient equilibrium sampling remains a major bottleneck.
Deep generative models have recently emerged as a promising alternative for equilibrium sampling (Noé et al., 2019; Albergo et al., 2019; Wirnsberger et al., 2020). A notable example is the Boltzmann Generator (BG) (Noé et al., 2019). By learning a diffeomorphic transformation between a simple prior (e.g., a Gaussian) and the complex molecular configuration space, BGs enable efficient proposal generation and exact likelihood evaluation. The tractable likelihood permits rigorous reweighting of generated configurations to the target Boltzmann distribution via importance sampling, yielding asymptotically exact equilibrium statistics. This enables amortized sampling (Tan et al., 2025b), where generation cost is substantially lower than MD simulations.
Figure 1: CG-BG workflow. (1) Training data are collected and mapped from atomistic configurations to CG beads. (2) A PMF network learns $U_{\eta}(\mathbf{R})$ from rapidly converged data, while a normalizing flow learns a proposal density $q_{\theta}(\mathbf{R})$. (3) CG samples from flow models are reweighted with the PMF to recover the target distribution $p(\mathbf{R})$ and compute unbiased thermodynamic observables.
In practice, however, scaling BGs to high-dimensional molecular systems remains challenging (Klein and Noé, 2024; Tan et al., 2025a). As system size increases, even expressive generative models exhibit diminishing overlap with the target distribution, resulting in high-variance importance weights and ineffective reweighting. Moreover, exact likelihood evaluation requires computing Jacobian determinants (Chen et al., 2018), which incurs significant computational overhead and scales poorly with dimensionality (Hutchinson, 1989). As a result, current BGs have mostly been applied to small systems, such as short peptides (Tan et al., 2025b; Rehman et al., 2025), often restricted to empirical implicit solvent models (Hawkins et al., 1995; Nguyen et al., 2013).
Coarse-graining (CG) provides a complementary approach by projecting atomistic configurations onto a lower-dimensional set of collective variables. This idea underlies Boltzmann Emulators (Lewis et al., 2025; Zheng et al., 2024; Jing et al., 2024; Zhu et al., 2026) and related generative surrogates (Daigavane et al., 2025; Schreiner et al., 2023; Plainer et al., 2025; Costa et al., 2025; dos Santos Costa et al., 2024; Diez et al., 2025; Vlachas et al., 2021; Xu et al., 2025), which reduce the number of degrees of freedom and allow models to effectively handle larger systems. However, these approaches face a circular dependency: training generally requires unbiased equilibrium data, which is precisely the computational bottleneck they aim to overcome. In most cases, CG surrogates are often trained on finite-time, non-equilibrium trajectories (Lewis et al., 2025; Zheng et al., 2024) and, unlike Boltzmann Generators, most lack a reweighting procedure to correct for this mismatch, resulting in biased statistics.
Present work. In this paper, we introduce Coarse-Grained Boltzmann Generators (CG-BGs, Fig. 1), a class of Boltzmann Generators that operate directly in CG coordinate  [^1] Our code is publicly available at https://github.com/tummfm/cg-bg.. CG-BGs combine generative modeling with importance sampling using a learned potential of mean force (PMF) as the target energy, enabling exact equilibrium sampling in a reduced-dimensional representation. This design provides a scalable pathway for sampling high-dimensional molecular systems. CG-BGs are particularly advantageous as they can be trained directly from rapidly converged data and effectively capture complex solvent-mediated effects.
Our main contributions are:
  • We introduce Coarse-Grained Boltzmann Generators, a scalable framework for equilibrium sampling in CG coordinate space. To our knowledge, this is the first BG formulation that incorporates machine learning potentials (MLPs) as the target energy for importance sampling.
  • We leverage enhanced sampling force matching to learn CG PMFs directly from rapidly converged simulations, eliminating reliance on unbiased equilibrium data and providing a correction mechanism for CG Boltzmann Emulators.
  • We demonstrate that CG-BGs capture solvent-mediated interactions in highly reduced representations, achieving improved accuracy over classical implicit solvent models while dramatically reducing computational cost relative to atomistic BGs.
2 Background and Preliminaries
Notation. We use lowercase variables for fine-grained (atomistic) quantities and uppercase variables for CG quantities.
We consider a many-body system with configuration $\mathbf{r}\in\mathbb{R}^{n}$ governed by a potential energy function $u(\mathbf{r})$. At thermodynamic equilibrium with temperature $T$, the system follows the Boltzmann distribution
$p(\mathbf{r})=\frac{e^{-\beta u(\mathbf{r})}}{Z},\quad Z=\int e^{-\beta u(\mathbf{r})}d\mathbf{r},$ (1)
where $\beta=(k_{B}T)^{-1}$ and $Z$ is the partition function. The goal of equilibrium sampling is to generate samples from $p(\mathbf{r})$ in order to compute observables $\mathbb{E}_{p}[\mathcal{O}]=\int\mathcal{O}(\mathbf{r})p(\mathbf{r})d\mathbf{r}$, and free energies. A dataset is considered converged if the empirical averages of observables statistically recover equilibrium expectations.
2.1 Boltzmann Generators and Emulators
Boltzmann Generators (BGs) (Noé et al., 2019) combine exact-likelihood generative models with importance sampling to obtain exact equilibrium statistics. Typically implemented using normalizing flows (Rezende and Mohamed, 2016), a BG defines a tractable proposal density $q_{\theta}(\mathbf{r})$ that approximates $p(\mathbf{r})$. Given samples $\mathbf{r}_{i}\sim q_{\theta}$, importance weights are assigned as
$w(\mathbf{r}_{i})=\frac{p(\mathbf{r}_{i})}{q_{\theta}(\mathbf{r}_{i})}\propto\frac{e^{-\beta u(\mathbf{r}_{i})}}{q_{\theta}(\mathbf{r}_{i})}.$ (2)
Unbiased estimates of equilibrium expectations are then obtained using the self-normalized importance sampling estimator
$\mathbb{E}_{p}[\mathcal{O}]\approx\frac{\sum_{i=1}^{N}w(\mathbf{r}_{i})\,\mathcal{O}(\mathbf{r}_{i})}{\sum_{i=1}^{N}w(\mathbf{r}_{i})}.$ (3)
Provided that $q_{\theta}$ overlaps sufficiently with $p$, this procedure converges to the exact Boltzmann average. This allows BGs to be trained on biased or non-equilibrium data, since reweighting restores equilibrium statistics at inference time.
Boltzmann Emulators adopt similar generative architectures but omit the reweighting step, estimating observables directly under $q_{\theta}$. While computationally efficient at inference, emulators are no longer asymptotically exact: their accuracy depends entirely on how closely $q_{\theta}$ approximates $p(\mathbf{r})$. This dependence necessitates high-quality unbiased training data, whose generation is often the primary computational bottleneck.
2.2 Continuous Normalizing Flows
Continuous normalizing flows (CNFs) extend discrete normalizing flows (Rezende and Mohamed, 2016; Dinh et al., 2014) by modeling density transformations as solutions to time-dependent ordinary differential equations (ODEs) (Chen et al., 2018). A vector field $v_{\theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, parameterized by a neural network, defines the dynamics
$\frac{d\mathbf{x}(t)}{dt}=v_{\theta}(t,\mathbf{x}(t)),\qquad\mathbf{x}(0)\sim p_{0},$ (4)
where $p_{0}$ is a simple prior distribution. The solution at time $t$ is given by
$\mathbf{x}(t)=\mathbf{x}(0)+\int_{0}^{t}v_{\theta}(\tau,\mathbf{x}(\tau))\,d\tau.$ (5)
The evolution of the log-density follows the instantaneous change-of-variables formula
$\log p_{t}(\mathbf{x}(t))=\log p_{0}(\mathbf{x}(0))-\int_{0}^{t}\nabla\cdot v_{\theta}(\tau,\mathbf{x}(\tau))\,d\tau,$ (6)
where $\nabla\cdot v_{\theta}$ denotes the divergence of the vector field.
While CNFs are commonly trained via maximum likelihood, Flow Matching (FM) (Albergo et al., 2023; Lipman et al., 2022; Liu et al., 2022) provides a simulation-free alternative. Conditional Flow Matching (CFM) (Tong et al., 2023) directly regresses the neural vector field $v_{\theta}$ onto a target conditional vector field $u_{t}(\mathbf{x}\mid z)$ that induces a prescribed probability path. The training objective is
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,z,\mathbf{x}\sim p_{t}(\cdot|z)}\big[\|v_{\theta}(t,\mathbf{x})-u_{t}(\mathbf{x}\mid z)\|^{2}\big],$ (7)
where $t\sim\mathcal{U}[0,1]$ and $z$ denotes a conditioning variable. A commonly used construction specifies the probability path through linear interpolation between paired samples (Albergo et al., 2023; Tong et al., 2023). Specifically, let $z=(\mathbf{x}_{0},\mathbf{x}_{1})$, where $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ are samples drawn from the source and target distributions, respectively. The interpolated state and corresponding vector field are
$\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1},\qquad u_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{1})=\mathbf{x}_{1}-\mathbf{x}_{0}.$ (8)
2.3 Coarse-Graining and Potentials of Mean Force
Coarse-graining reduces dimensionality by mapping atomistic configurations $\mathbf{r}\in\mathbb{R}^{n}$ to a lower-dimensional set of collective variables (CVs), or beads, $\mathbf{R}\in\mathbb{R}^{N}$ ($N\ll n$), typically selected to capture the slow degrees of freedom, via a mapping $\mathbf{R}=\Xi(\mathbf{r})$. From a bottom-up perspective (Noid et al., 2008; Jin et al., 2022), the goal of coarse-graining is to define a suitable set of CG variables and parameterize an effective potential such that the CG model reproduces the marginal equilibrium distribution of the underlying atomistic system:
$p(\mathbf{R})=\int p(\mathbf{r})\,\delta(\Xi(\mathbf{r})-\mathbf{R})\,d\mathbf{r}.$ (9)
This marginal distribution admits a Boltzmann form,
$p(\mathbf{R})\propto{e^{-\beta U(\mathbf{R})}},$ (10)
where the effective energy $U(\mathbf{R})$, known as the potential of mean force, is defined up to an additive constant as
$U(\mathbf{R})=-k_{B}T\ln\int e^{-\beta u(\mathbf{r})}\,\delta(\Xi(\mathbf{r})-\mathbf{R})\,d\mathbf{r}.$ (11)
The PMF captures both energetic and entropic contributions from the eliminated degrees of freedom and generally induces many-body, state-dependent interactions (Krishna et al., 2009).
In practice, the PMF is intractable due to the high-dimensional integral over $\mathbf{r}$. Classical CG force fields (Marrink et al., 2007; Souza et al., 2021) approximate $U(\mathbf{R})$ using fixed functional forms, which may lack sufficient expressivity. Recent approaches parameterize $U(\mathbf{R})$ using neural networks trained via variational principles such as force matching (Noid et al., 2008; Wang et al., 2019) or relative entropy minimization (Shell, 2008; Thaler et al., 2022).
3 Coarse-Grained Boltzmann Generators
Atomistic BGs provide exact equilibrium sampling but scale poorly to large systems, while CG Boltzmann Emulators improve scalability at the cost of persistent bias due to the absence of reweighting.
We propose CG-BGs that reconcile scalability with statistical rigor by performing generative modeling and importance sampling directly in CG coordinate space. Rather than targeting the full atomistic distribution, CG-BGs target the marginal distribution $p(\mathbf{R})$ defined by the PMF.
CG-BGs have two major components: a flow-based generative model to propose CG samples (similar to atomistic BGs) and a machine learning PMF acting as an energy predictor for importance sampling. Unlike atomistic MLPs trained on potential energy surfaces (Behler and Parrinello, 2007; Blank et al., 1995), CG MLPs have no ground-truth energy labels, since the PMF is a free energy that includes entropic contributions from eliminated degrees of freedom. We next describe how the PMF is learned from atomistic simulations and outline the CG-BG workflow (Fig. 1).
3.1 Variational Force Matching
Variational Force Matching (VFM) (Noid et al., 2008), also known as multiscale coarse-graining, is a bottom-up approach for learning the PMF from atomistic forces.
The central premise is that the CG forces predicted by the model should match, on average, the instantaneous atomistic forces projected onto the CG coordinates, denoted as $\mathcal{F}_{\mathrm{proj}}(\mathbf{r})$. The gradient of the exact PMF satisfies:
$\nabla U(\mathbf{R})=\mathbb{E}_{p(\mathbf{r}\mid\mathbf{R})}\big[\mathcal{F}_{\mathrm{proj}}(\mathbf{r})\big].$ (12)
where the expectation is taken over the fiber distribution (Hummerich et al., 2025), i.e., the conditional probability of a fine-grained configuration $\mathbf{r}$ given its CG representative $\mathbf{R}$.
From a learning perspective, instantaneous projected forces provide noisy samples to approximate the target mean force. Given a dataset $\mathcal{D}$ of atomistic configurations, a parameterized CG potential $U_{\eta}(\mathbf{R})$ is trained by minimizing
$\mathcal{L}_{\mathrm{VFM}}(\eta)=\mathbb{E}_{\mathbf{r}\sim\mathcal{D}}\Big[\big\|\nabla_{\mathbf{R}}U_{\eta}(\Xi(\mathbf{r}))+\mathcal{F}_{\mathrm{proj}}(\mathbf{r})\big\|_{2}^{2}\Big].$ (13)
When $\mathcal{D}$ is sampled from equilibrium, this objective minimizes the Fisher divergence between the model distribution $p_{\eta}(\mathbf{R})$ and the true marginal $p(\mathbf{R})$. Theoretical error bounds follow from Log-Sobolev inequalities (Proof in §A.1):
{restatable} propositionpropcom Let $p^{*}(\mathbf{R})\propto e^{-\beta U^{*}(\mathbf{R})}$ be the true marginal and $p_{\eta}(\mathbf{R})\propto e^{-\beta U_{\eta}(\mathbf{R})}$ the learned distribution. If $p^{*}$ satisfies a Logarithmic Sobolev Inequality (LSI) with constant $\rho>0$. Then, the Kullback-Leibler divergence between the learned and true distributions is bounded by the expected squared force error: $\mathcal{D}_{\mathrm{KL}}(p_{\eta}\|p^{*})\leq\frac{\beta^{2}}{2\rho}\mathbb{E}_{p_{\eta}}\left[\|\nabla U_{\eta}(\mathbf{R})-\nabla U^{*}(\mathbf{R})\|^{2}\right].$ (14)
While global LSI conditions are strong assumptions for multimodal PMFs (Vempala and Wibisono, 2019), this proposition theoretically motivates minimizing force errors as a proxy for distributional accuracy. This is important for the subsequent importance sampling step, as the quality of the reweighted ensemble depends directly on the divergence between the learned target and the true marginal.
3.2 Enhanced Sampling for Force Matching
Standard force matching relies on unbiased equilibrium data, which can be prohibitively expensive to obtain for systems with metastable states and large free-energy barriers—a limitation shared by Boltzmann Emulators. Moreover, the Boltzmann distribution decays exponentially with energy, leading to sparse sampling of high-energy transition regions that are critical for accurately reconstructing the global PMF.
Enhanced sampling force matching (ESFM) (Chen et al., 2026) overcomes these limitations by exploiting the invariance of the fiber (conditional) distribution under coarse-grained biasing.
{restatable} propositionfiber (Chen et al. (2026)) Let $V(\mathbf{R})$ be a bias potential depending only on the coarse-grained coordinates. The conditional distribution of atomistic configurations given $\mathbf{R}$ remains unchanged: $p_{V}(\mathbf{r}\mid\mathbf{R})=p(\mathbf{r}\mid\mathbf{R}).$ (15)
Since the mean force $\nabla U(\mathbf{R})$ is the expectation of the projected forces over this conditional distribution (Eq. 12), the force regression targets remain unbiased even when sampled from a biased equilibrium ensemble. ESFM therefore minimizes
$\mathcal{L}_{\mathrm{ESFM}}(\eta)=\mathbb{E}_{\mathbf{r}\sim\mathcal{D}_{\mathrm{bias}}}\Big[\big\|\nabla_{\mathbf{R}}U_{\eta}(\Xi(\mathbf{r}))+\mathcal{F}_{\mathrm{proj}}(\mathbf{r})\big\|_{2}^{2}\Big],$ (16)
where $\mathcal{D}_{\mathrm{bias}}$ is a rapidly converged dataset generated via enhanced sampling and $\mathcal{F}_{\mathrm{proj}}(\mathbf{r})$ denotes forces recomputed using the unbiased atomistic potential.
{restatable} propositionesfm (Chen et al. (2026)) Minimizing $\mathcal{L}_{\mathrm{ESFM}}$ yields the same global optimum as standard force matching loss $\mathcal{L}_{\mathrm{VFM}}$, assuming sufficient model expressivity.
Together, Propositions 3.2 (Proof in §A.2) and 3.2 establish that ESFM enables accurate PMF learning from rapidly converged data, benefiting from faster convergence and improved coverage of transition regions.
3.3 The CG-BG Workflow
After training, the learned PMF $U_{\eta}$ defines the target energy for importance sampling, rather than serving as a force field for MD integration. Let $q_{\theta}(\mathbf{R})$ denote the density induced by the trained flow model. Importance weights are computed as
$w(\mathbf{R})\propto\frac{\exp(-\beta U_{\eta}(\mathbf{R}))}{q_{\theta}(\mathbf{R})}.$ (17)
Provided $U_{\eta}$ accurately approximates the true PMF on the support of $q_{\theta}$, this procedure gives unbiased estimates under $p(\mathbf{R})$.
The reliability of importance reweighting is quantified by the effective sample size (ESS) (Kish, 1965),
$\mathrm{ESS}=\frac{\left(\sum_{i=1}^{B}w(\mathbf{R}_{i})\right)^{2}}{\sum_{i=1}^{B}w(\mathbf{R}_{i})^{2}}.$ (18)
The ESS depends critically on the overlap between the generative density $q_{\theta}$ and the target density defined by $U_{\eta}$. In practice, machine learning potentials may exhibit unphysical extrapolation outside the training domain, and generative models may occasionally produce high-energy artifacts, both of which can lead to weight degeneracy. To improve robustness, we apply a weight clipping strategy (Tan et al., 2025a; Gloy and Olsson, 2025; Moqvist et al., 2025) to truncate statistical outliers before computing expectations (See §G.2). The complete CG-BG training and sampling pipeline is summarized in Fig. 1.
4 Experiments
We evaluate CG-BGs on the Müller–Brown (MB) potential and alanine dipeptide (capped alanine). Additional experimental details including architectures are provided in §C and §D. CG-BG samples are generated and reweighted following the algorithms described in §F.
Datasets. For both systems, we construct unbiased and biased datasets for training and evaluation. Biased datasets are generated using enhanced sampling methods to accelerate exploration of configurational space. For the MB system, datasets are generated via Langevin dynamics (§B.1), with umbrella sampling (Torrie and Valleau, 1977) applied in the biased setting to facilitate transitions between metastable basins. For alanine dipeptide, we generate datasets using both explicit and implicit solvent models (§B.2). Explicit solvent datasets, used for training, are produced using a classical force field (Lindorff-Larsen et al., 2010) and include long unbiased trajectories as well as biased simulations from well-tempered metadynamics (Barducci et al., 2008). Implicit solvent datasets are generated with the same force field combined with a generalized Born solvent model, using two parameterizations (OBC1 and OBC2) (Onufriev et al., 2004). The explicit solvent trajectories are further coarse-grained using either a Heavy Atom mapping (Fig. 3a), which retains all heavy atoms, or a Core Beta mapping (Fig. 4a), which preserves the backbone and $C_{\beta}$ atoms. Full details on dataset generation and enhanced sampling procedures are provided in §B.
Baselines. Unlike previous BG work (Klein and Noé, 2024; Tan et al., 2025a, b), which often treats implicit solvent simulations as ground truth, we use explicit solvent simulations as reference and regard implicit solvent models as baselines. For alanine dipeptide, CG reduces the effective dimensionality to either 10 beads (Heavy Atom) or 6 beads (Core Beta). To quantify the computational gain via coarse-graining, we use the full 22-atom system as a baseline.
Metrics. We report ESS, Jenson-Shannon (JS) divergence, and PMF error. JS divergence is computed between the sampled and reference dihedral angle free energy profiles. The PMF error is defined as the squared distance between the negative logarithms of the sampled and reference densities, which gives more weight to low-density regions compared to JS divergence (Plainer et al., 2025; Durumeric et al., 2024). Energy histograms and free energy profiles of $\phi$ dihedral are shown in the main text, while additional results, including $\psi$ dihedral free energy profiles, Ramachandran plots (Ramachandran, 1963), bond length distributions, and weight clipping ablations, are provided in §G.
4.1 Recovering Equilibrium Distributions from Biased and Unbiased Data
We first demonstrate that CG-BGs inherit the importance reweighting capability of atomistic BGs, enabling recovery of unbiased equilibrium statistics from flow models trained on either biased or unbiased trajectories.
For the MB system (Fig. 2a–b), coarse-graining corresponds to projection onto the $x$-coordinate, yielding an analytically exact reference $U(x)=-k_{B}T\ln\int\exp\!\big(-\beta u(x,y)\big)\,dy$ (Fig. 2c–d). For alanine dipeptide, we use the Heavy Atom mapping, which removes hydrogens and solvent degrees of freedom (Fig. 3a). Throughout §4.1 and §4.2, the PMF is learned from rapidly converged datasets using ESFM (§3.2).
Figure 2: CG-BGs on the MB potential. (a) Two-dimensional MB potential energy surface (functional form in §B). (b) Marginal probability density along the $x$ coordinate. (c) Free energy profiles before and after reweighting for CG-BGs, where flow is trained on unbiased data, compared with the exact solution and MD reference. (d) Same as (c), but for flow trained on biased data.
As shown in Fig. 2c and Fig. 3c, despite training on long unbiased datasets and using highly expressive models, the raw flow proposals deviate from the MD reference. These discrepancies arise from low-quality samples generated by the flow model, a limitation intrinsic to Boltzmann Emulators that cannot be systematically corrected without reweighting. After reweighting, however, CG-BGs successfully recover equilibrium free energy profiles in close agreement with MD references (Fig. 2c and Fig. 3b-c). The reweighted distributions accurately reproduce relative basin populations and capture transition regions. This is further illustrated by the Ramachandran plots (Fig. 7). Unreweighted proposals exhibit noisy samples across transition regions and metastable basins (Fig. 7b), which are assigned low importance weights and effectively filtered out after reweighting (Fig. 7c). Quantitative metrics in Tab. 2 confirm close agreement between reweighted samples and target equilibrium ensemble.
We further evaluate CG-BGs trained on biased or short non-equilibrium trajectories (umbrella sampling for MB and 10 ns WT-MetaD with $\gamma=1.5$ for alanine dipeptide), reflecting realistic settings where long unbiased simulations are infeasible. As shown in Fig. 2d, Fig. 3d and Fig. 7c, while the flow proposals exhibit larger deviations from the MD reference, importance reweighting consistently recovers accurate equilibrium statistics (Tab. 2).
Figure 3: CG-BGs on alanine dipeptide (Heavy Atom). (a) Heavy Atom mapping. (b) Potential energy distributions under the learned PMF before and after reweighting, compared with the MD reference. (c) $\phi$ dihedral free energy profile before and after reweighting for CG-BGs, where flow is trained on 500 ns unbiased data, alongside the MD reference. (d) Same as (c), but for flow trained on a 10 ns WT-MetaD ($\gamma=1.5$) dataset.
Notably, after reweighting, CG-BGs outperform implicit solvent MD baselines (Tab. 2), highlighting the advantage of learning PMFs from explicit solvent simulations. While implicit solvent models perform reasonably well for alanine dipeptide, prior work has reported drastic discrepancies for more complex systems (Chen et al., 2021). This constitutes an improvement over previous BG approaches, which rely on implicit solvent models for reweighting and are therefore fundamentally limited by solvent approximation error. In other words, such models can at best achieve the accuracy of implicit solvent baselines (Tab. 2). These results show that CG-BGs generate CG equilibrium samples consistent with the target distribution, without requiring long unbiased MD simulations, and offer a practical route to correct systematic biases in existing Boltzmann Emulators.
4.2 Effect of Coarse-Graining Resolution on Accuracy and Efficiency
We next examine how the choice of CG resolution affects both modeling accuracy and computational efficiency. To this end, we consider a coarser representation of alanine dipeptide, the Core Beta mapping (Fig. 4a).
Despite the reduced resolution, CG-BGs trained with the Core Beta mapping remain capable of recovering equilibrium statistics after reweighting (Fig. 4c–d and Fig. 10), with quantitative metrics reported in Tab. 2. While performance remains strong for alanine dipeptide, coarser mappings are known to induce a gradual loss of accuracy for more complex systems (Görlich and Zavadlav, 2025). This degradation arises from increased degeneracy in the CG representation. For a given CG coordinate $\mathbf{R}$, many atomistic microstates $\mathbf{r}_{i}$} satisfy $\Xi(\mathbf{r}_{i})=\mathbf{R}$ while exerting different projected forces $\mathcal{F}_{\mathrm{proj}}(\mathbf{r}_{i})$. As a result, the conditional mean force $\mathbb{E}_{p(\mathbf{r}\mid\mathbf{R})}\!\left[\mathcal{F}_{\mathrm{proj}}(\mathbf{r})\right]$ exhibits higher variance under coarser mappings, introducing irreducible uncertainty and making the force matching problem intrinsically more challenging.
Figure 4: CG-BGs on alanine dipeptide (Core Beta). (a) Core Beta mapping. (b) Potential energy distributions under the learned PMF before and after reweighting, compared with the MD reference. (c) $\phi$ dihedral free energy profile before and after reweighting for CG-BGs, where flow is trained on 500 ns unbiased data, alongside the MD reference. (d) Same as (c), but for flow trained on a 10 ns WT-MetaD ($\gamma=1.5$) dataset .
We further benchmark the computational cost of CG-BGs across different resolutions (Tab. 1). As expected, the Core Beta mapping substantially reduces both training and inference time compared to the Heavy Atom and full atomistic representations, with particularly large gains at inference due to cheaper Jacobian evaluations. Note that the reported all atom baseline generates only solute coordinates; generating full configurations with explicit solvent, which would be the proper reference, is computationally infeasible. Overall, these results highlight the favorable accuracy–efficiency trade-off of CG-BGs. Additional comparisons with prior work are provided in Tab. 4.
Table 1: Training and inference time across different CG mappings. Inference times correspond to $10^{4}$ generated samples. All Atom denotes generating the full solute configuration without solvent.
Stage Core Beta Heavy Atom All Atom
Training 0.83h 1.39h 3.93h
Inference 0.84min 3.46min 13.92min
Total 0.84h 1.45h 4.16h
Table 2: Quantitative comparison for alanine dipeptide across CG resolutions and baseline models. CG-BGs results are reported after reweighting; Biased denotes flow trained on WT-MetaD datasets, while $\text{PMF}_{U}$ indicates PMFs learned from long unbiased MD data.
Model JS ($\downarrow$) PMF ($\downarrow$) ESS ($\uparrow$)
CG-BGs (Reweighted)
Heavy Atom 0.0075 $\pm$ 0.0001 $0.2493\pm 0.0126$ $0.2325\pm 0.0006$
Heavy Atom (Biased) $0.0090\pm 0.0002$ $0.2758\pm 0.0109$ $0.1936\pm 0.0005$
Core Beta $0.0079\pm 0.0001$ $0.2519\pm 0.0136$ 0.2510 $\pm$ 0.0006
Core Beta (Biased) $0.0081\pm 0.0001$ 0.2325 $\pm$ 0.0093 $0.2239\pm 0.0005$
Heavy Atom ($\text{PMF}_{U}$) $0.0218\pm 0.0002$ $0.3611\pm 0.0138$ $0.2222\pm 0.0006$
Implicit Solvent Baselines (Generalized Born Models)
GB (OBC1) $0.0157\pm 0.0002$ $0.3709\pm 0.0092$
GB (OBC2) $0.0182\pm 0.0002$ $0.4028\pm 0.0095$
4.3 Simulation-Free Evaluation of Learned PMFs
Beyond equilibrium sampling, CG-BGs also provide a practical tool for evaluating learned PMFs without running MD. The flow model generates CG configurations from a proposal distribution approximating the equilibrium ensemble, and importance reweighting maps these samples to the Boltzmann distribution induced by a given PMF. As a result, equilibrium observables for any candidate PMF can be estimated using a single set of generated samples and associated weights, without additional simulations. In contrast to traditional validation pipelines that require long MD runs for each model, CG-BGs enable rapid, one-shot benchmarking of CG potentials.
We leverage this capability to benchmark PMFs trained under different data regimes, comparing models learned from long unbiased MD trajectories ($\text{PMF}_{U}$) with those trained on a rapidly converged biased dataset ($\text{PMF}_{B}$). Quantitative metrics (Tab. 2) computed from reweighted samples provide immediate feedback on model accuracy, while visual diagnostics including probability densities(Fig. 5a), dihedral free energy profiles (Fig. 5b), and Ramachandran plots (Fig. 9) qualitative validation against atomistic references. Consistent with prior observations (Chen et al., 2026; Görlich and Zavadlav, 2025), $\text{PMF}_{U}$ fails to capture the correct metastable population of the $\phi$ angle, whereas $\text{PMF}_{B}$ shows improved agreement.
While demonstrated here for CG models, this simulation-free evaluation paradigm naturally extends to atomistic MLPs. CG-BGs therefore serve not only as efficient equilibrium sampling methods, but also as diagnostic tools for rapid comparison and validation of learned energy models.
Figure 5: Simulation-free benchmarking of learned CG PMFs using CG-BGs (Heavy Atom). (a) Probability density of the $\phi$ dihedral angle after reweighting with PMFs trained on unbiased ($\text{PMF}_{U}$) and rapidly converged biased datasets ($\text{PMF}_{B}$), compared with the MD reference and flow proposal (trained on unbiased data). (b) Corresponding $\phi$ dihedral free energy profiles after reweighting, alongside the MD reference and flow proposal.
5 Related Work
Generative models for equilibrium molecular sampling. Generative models have emerged as powerful tools for sampling Boltzmann distributions (Olsson, 2026; Klein et al., 2023; Rotskoff, 2024; Aranganathan et al., 2025; Janson and Feig, 2025). In particular, Boltzmann Generators (Noé et al., 2019) represent an influential class of such models. Subsequent work has improved BGs in various ways (von Klitzing et al., 2025; Schebek and Rogal, 2025), including incorporating inductive biases (Köhler et al., 2020) and enhancing transferability across chemical and thermodynamic conditions (Klein and Noé, 2024; Dibak et al., 2022; Moqvist et al., 2025; Invernizzi et al., 2022). Despite efforts to employ scalable architectures (Tan et al., 2025a; Zhai et al., 2024) or reduce the cost of likelihood evaluation (Gloy and Olsson, 2025; Rehman et al., 2025; Peng and Gao, 2025), atomistic BGs remain limited in scalability. CG Boltzmann Emulators (Lewis et al., 2025; Zheng et al., 2024) bypass the reweighting step, enabling larger-scale applications at the cost of asymptotic exactness. To bridge the resolution gap, backmapping (Chennakesavalu et al., 2023; Hummerich et al., 2025; Wang et al., 2022) or other reconstruction strategies (Tamagnone et al., 2024; Schopmans and Friederich, 2024; Stupp and Koutsourelakis, 2025) are often employed to recover atomistic detail. Another active line of research considers neural samplers for sampling unnormalized densities (Akhound-Sadegh et al., 2024; Midgley et al., 2023; He et al., 2025; Potaptchik et al., 2025; Liu et al., 2025), which also have applications in molecular systems (Nam et al., 2025; Havens et al., 2025).
Coarse-grained machine learning potentials. CG potentials have historically been parameterized using either top-down or bottom-up approaches. In top-down methods, model parameters are adjusted to reproduce macroscopic observables, such as experimental measurements (Marrink et al., 2007; Thaler and Zavadlav, 2021; Fuchs and Zavadlav, 2025). In contrast, bottom-up approaches aim to construct CG models that reproduce the marginal equilibrium distribution or free energy landscape of the underlying atomistic system (Jin et al., 2022; Noid, 2013). Prominent bottom-up methods include variational force matching (Noid et al., 2008) and relative entropy minimization (Shell, 2008). Recent progress in machine learning has opened new opportunities for learning CG MLPs (Plainer et al., 2025; Kohler et al., 2023; Arts et al., 2023; Wang et al., 2019; Zhang et al., 2018). The development of CG MLPs can be viewed as an extension of broader effort to construct accurate machine learning interatomic potentials from first-principles calculations. (Batatia et al., 2022; Wood et al., 2025; Batzner et al., 2022; Unke et al., 2021; Kabylda et al., 2025). Despite these successes, learning transferable and efficient CG MLPs remains an open challenge (Charron et al., 2025; Mirarchi et al., 2024; Majewski et al., 2023; Durumeric et al., 2023).
6 Conclusion
This work introduces CG-BGs, a scalable framework for equilibrium sampling of coarse-grained molecular systems. By targeting the marginal equilibrium distribution defined by the PMF, CG-BGs reduce the effective dimensionality of the sampling problem while preserving thermodynamic consistency through importance reweighting. The underlying PMF can be learned from rapidly converged data using enhanced sampling force matching, providing a correction mechanism for existing CG Boltzmann Emulators. Even at high levels of coarse-graining, CG-BGs capture solvent-mediated and many-body effects, and enable one-shot, simulation-free evaluation of CG MLPs.
Limitations. The current approach relies on predefined collective variables for coarse-graining and enhanced sampling, which may be nontrivial to identify for complex systems. Recent advances in collective variable discovery (Zhang et al., 2024; Ribeiro et al., 2018; Chen and Ferguson, 2018; Herringer et al., 2023; Mehdi et al., 2024) and uncertainty quantification (Zaverkin et al., 2024; Musil et al., 2019) provide promising avenues to address these challenges.
Future work. Extending CG-BGs to larger, more complex systems is a natural next step, leveraging demonstrated transferability of flow-based generative models and MLPs. The simulation-free evaluation paradigm introduced here also opens new opportunities, including integrating CG-BGs with experimental data by incorporating it into the learned PMF, and accelerating differentiable training strategies (Thaler and Zavadlav, 2021; Han and Yu, 2025) by replacing iterative simulations with high-quality generated samples.
Broader Impact
This work focuses on sampling from Boltzmann distributions, a problem of broad interest in machine learning and AI for Science, with applications in both statistical physics and molecular simulations. We introduce Coarse-Grained Boltzmann Generators, which can be trained on molecular systems and applied to tasks such as drug and material discovery. While we do not anticipate immediate negative impacts, we encourage careful consideration when scaling these methods to prevent potential misuse.
Acknowledgements
We thank Simon Olsson, Franz Görlich, Nuno Costa, and Paul Fuchs for fruitful discussions and helpful feedback. This work was funded by the European Union through the ERC (StG SupraModel) - 101077842. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Proofs
A.1 Proof of Proposition 3.1
[…]
Proof.
The Kullback-Leibler divergence is defined as:
$\mathcal{D}_{\mathrm{KL}}(p_{\eta}\|p^{*})=\int p_{\eta}(\textbf{R})\log\frac{p_{\eta}(\mathbf{R})}{p^{*}(\mathbf{R})}d\mathbf{R}.$ (19)
The Fisher Divergence (or relative Fisher information) between $p_{\eta}$ and $p^{*}$ is defined as:
$\mathcal{J}(p_{\eta}\|p^{*})=\int p_{\eta}(\mathbf{R})\left\|\nabla\log p_{\eta}(\mathbf{R})-\nabla\log p^{*}(\mathbf{R})\right\|^{2}d\mathbf{R}.$ (20)
Since $p_{\eta}(\mathbf{R})=Z_{\eta}^{-1}e^{-\beta U_{\eta}(\mathbf{R})}$ and $p^{*}(\mathbf{R})=(Z^{*})^{-1}e^{-\beta U^{*}(\mathbf{R})}$, the gradients of the log-densities are directly proportional to the forces:
$\nabla\log p_{\eta}(\mathbf{R})=-\beta\nabla U_{\eta}(\mathbf{R}),\quad\nabla\log p^{*}(\mathbf{R})=-\beta\nabla U^{*}(\mathbf{R}).$ (21)
Substituting these into the definition of the Fisher Divergence:
$\mathcal{J}(p_{\eta}\|p^{*})$ $=\int p_{\eta}(\mathbf{R})\left\|(-\beta\nabla U_{\eta}(\mathbf{R}))-(-\beta\nabla U^{*}(\mathbf{R}))\right\|^{2}d\mathbf{R}$ (22)
$=\beta^{2}\int p_{\eta}(\mathbf{R})\left\|\nabla U_{\eta}(\mathbf{R})-\nabla U^{*}(\mathbf{R})\right\|^{2}d\mathbf{R}$ (23)
$=\beta^{2}\mathbb{E}_{p_{\eta}}\left[\|\mathcal{F}_{\eta}(\mathbf{R})-\mathcal{F}^{*}(\mathbf{R})\|^{2}\right].$ (24)
We assume that the target distribution $p^{*}$ satisfies a Logarithmic Sobolev Inequality (LSI) with constant $\rho>0$. By definition, this inequality implies that for any distribution $p_{\eta}$ absolutely continuous with respect to $p^{*}$:
$\mathcal{D}_{\mathrm{KL}}(p_{\eta}\|p^{*})\leq\frac{1}{2\rho}\mathcal{J}(p_{\eta}\|p^{*}).$ (25)
Substituting our expression for the Fisher Divergence into the LSI yields the final bound:
$\mathcal{D}_{\mathrm{KL}}(p_{\eta}\|p^{*})\leq\frac{1}{2\rho}\left(\beta^{2}\mathbb{E}_{p_{\eta}}\left[\|\nabla U_{\eta}(\mathbf{R})-\nabla U^{*}(\mathbf{R})\|^{2}\right]\right).$ (26)
Dividing out the constants concludes the proof. ∎
A.2 Proof of Proposition 3.2
[…]
Proof.
Let $p(\mathbf{r})=Z^{-1}e^{-\beta u(\mathbf{r})}$ be the unbiased equilibrium distribution. By definition, the unbiased marginal distribution is
$p(\mathbf{R})=\int p(\mathbf{r})\delta(\Xi(\mathbf{r})-\mathbf{R})d\mathbf{r}=\frac{1}{Z}\int e^{-\beta u(\mathbf{r})}\delta(\Xi(\mathbf{r})-\mathbf{R})d\mathbf{r}.$ (27)
Rearranging this yields the identity for the unnormalized marginal:
$\int e^{-\beta u(\mathbf{r})}\delta(\Xi(\mathbf{r})-\mathbf{R})d\mathbf{r}=Zp(\mathbf{R}).$ (28)
Now, consider the biased distribution $p_{V}(\mathbf{r})=Z_{V}^{-1}e^{-\beta(u(\mathbf{r})+V(\Xi(\mathbf{r})))}$. The biased marginal distribution is:
$p_{V}(\mathbf{R})$ $=\int\frac{e^{-\beta u(\mathbf{r})}e^{-\beta V(\Xi(\mathbf{r}))}}{Z_{V}}\delta(\Xi(\mathbf{r})-\mathbf{R})d\mathbf{r}$ (29)
[…]
$=\frac{Z}{Z_{V}}e^{-\beta V(\mathbf{R})}p(\mathbf{R}).$ (31)
Finally, substituting this into the definition of the conditional distribution:
$p_{V}(\mathbf{r}\mid\mathbf{R})=\frac{p_{V}(\mathbf{r})\delta(\Xi(\mathbf{r})-\mathbf{R})}{p_{V}(\mathbf{R})}=\frac{Z_{V}^{-1}e^{-\beta u(\mathbf{r})}e^{-\beta V(\mathbf{R})}\delta(\dots)}{\frac{Z}{Z_{V}}e^{-\beta V(\mathbf{R})}p(\mathbf{R})}=\frac{e^{-\beta u(\mathbf{r})}\delta(\dots)}{Zp(\mathbf{R})}=p(\mathbf{r}\mid\mathbf{R}).$ (32)
Appendix B Datasets
B.1 Müller-Brown Potential
Potential Parameters. For the two-dimensional toy system, we use a Müller–Brown potential defined as
$u(x,y)=u_{1}(x,y)+u_{2}(x,y)+u_{3}(x,y)+u_{4}(x,y),$ (33)
with (Raja et al., 2025)
$u_{1}(x,y)$ $=-17.3\,\exp\left[-0.0039(x-48)^{2}-0.0391(y-8)^{2}\right],$
$u_{2}(x,y)$ $=-8.7\,\exp\left[-0.0039(x-32)^{2}-0.0391(y-16)^{2}\right],$
$u_{3}(x,y)$ $=-14.7\,\exp\left[-0.0254(x-24)^{2}+0.043(x-24)(y-32)-0.0254(y-32)^{2}\right],$
$u_{4}(x,y)$ =  $=\;\;1.3\,\exp\left[0.00273(x-16)^{2}+0.0023(x-16)(y-24)+0.00273(y-24)^{2}\right].$
Umbrella Sampling. For umbrella sampling, we introduce a biasing potential along the $x$ coordinate,
$V_{x}(x)=-4\,\exp\left[-\frac{(x-32.0)^{2}}{2\cdot 5^{2}}\right].$
This enables better sampling of the configurations around $\mathbf{x}_{0}=32$, allowing a better representation of transition regions that are otherwise rarely visited in unbiased trajectories.
Simulation Details. For the MB dataset generation, we perform two-dimensional Langevin dynamics with a time step of $0.1$, mass $m=1.0$, friction coefficient $\gamma=0.1$, and temperature $k_{\mathrm{B}}T=1.0$. Ten independent trajectories of length $10^{7}$ steps are generated, with initial positions sampled uniformly from $10,50]^{2}$ and initial velocities drawn from a Gaussian distribution with standard deviation $0.1$. Configurations are recorded every 10 steps. The dynamics follow
$m\ddot{\mathbf{r}}=-\nabla(u(\mathbf{r})+V(x))-\gamma m\dot{\mathbf{r}}+\sqrt{2\gamma k_{\mathrm{B}}Tm}\,\boldsymbol{\eta}(t),$ (34)
where $V(x)$ is the applied bias, and $\boldsymbol{\eta}(t)$ denotes Gaussian white noise. For each saved configuration, unbiased forces from $\nabla u$ are computed and stored.
B.2 Alanine Dipeptide
Force Fields. For the explicit solvent dataset, alanine dipeptide is parameterized using the AMBER99SB-ILDN force field (Lindorff-Larsen et al., 2010) and solvated in a cubic box of TIP3P water molecules. All explicit solvent simulations are carried out with GROMACS (Van Der Spoel et al., 2005). For the implicit solvent dataset, the same force field is used together with the generalized Born (OBC1/OBC2) implicit solvent model, and simulations are performed using OpenMM (Eastman et al., 2023).
Well-Tempered Metadynamics. We perform well-tempered metadynamics (WT-MetaD) (Barducci et al., 2008) simulations of alanine dipeptide in explicit solvent using GROMACS coupled with PLUMED (Bonomi et al., 2009). The backbone dihedral angles $\phi$ (C–N–Cα–C) and $\psi$ (N–Cα–C–N) are chosen as collective variables. Gaussian hills with height $1.2\,\text{kJ/mol}$ and width $0.35\,\text{rad}$ are deposited every 500 integration steps ($1\,\text{ps}$). Datasets are generated with bias factors $\gamma=1.5$ and $\gamma=9$, where $\gamma=1.5$ is used to train flow model for biased proposal generation and $\gamma=9$ is used for enhanced sampling force matching. Each simulation is run for $2\times 10^{7}$ steps ($10\,\text{ns}$) with a time step of $0.5\,\text{fs}$ and no bond constraints. Positions and forces are recorded every 40 steps ($0.02\,\text{ps}$), yielding $5\times 10^{5}$ saved configurations per dataset. To ensure unbiased force labels, all forces are recomputed by rerunning the saved trajectories in GROMACS without the metadynamics bias using the mdrun -rerun functionality. This guarantees that each configuration is associated with forces from the underlying unbiased potential.
Simulation Details. All simulations are performed in the NVT ensemble at a temperature of 300 K. For explicit solvent systems, alanine dipeptide is solvated in a cubic box of TIP3P water with a side length of 3.7 nm. After energy minimization, production dynamics are carried out using a velocity-rescale thermostat (time constant 0.1 ps). Long-range electrostatics are treated using the particle mesh Ewald method, and van der Waals interactions are truncated at 1.0 nm. We summarize the dataset configurations used in this work in Tab. 3.
Table 3: Overview of simulation details for alanine dipeptide datasets.
Solvent Dataset Method Length Timestep Constraints
Explicit (TIP3P) Unbiased MD None 500 ns 0.5 fs None
Explicit (TIP3P) Biased MD (for CNF) WT-MetaD ($\gamma=1.5$) 10 ns 0.5 fs None
Explicit (TIP3P) Biased MD (for PMF) WT-MetaD ($\gamma=9$) 10 ns 0.5 fs None
Implicit (OBC1) Unbiased MD None 500 ns 2 fs H-bonds
Implicit (OBC2) Unbiased MD None 500 ns 2 fs H-bonds
Appendix C Experimental Details for Conditional Normalizing Flows
C.1 Architecture
Müller-Brown Potential. For MB potential, we use a multilayer perceptron augmented with time conditioning for flow matching. The network consists of three hidden layers with width 96. Flow time is embedded into a 16-dimensional time embedding and concatenated with the input.
Alanine Dipeptide. For the alanine dipeptide, we use an adapted Graph Transformer architecture (Plainer et al., 2025; Arts et al., 2023; Shi et al., 2020). Given bead positions $\mathbf{x}_{i}$ and bead features $\mathbf{h}_{i}$, edge attributes are constructed as
$\mathbf{d}_{ij}=\mathbf{x}_{i}-\mathbf{x}_{j},\quad r_{ij}=\lVert\mathbf{d}_{ij}\rVert,\quad\mathbf{e}_{ij}=[\mathbf{d}_{ij},r_{ij}],$ (35)
and node attributes are initialized as
$\mathbf{n}^{(0)}_{i}=[\mathbf{h}_{i},\mathbf{x}_{i},t],$ (36)
where $t$ denotes the flow time. The bead features and time are embedded into 16- and 4-dimensional vectors, respectively, concatenated with positions, and projected to 128-dimensional node embeddings via a linear layer. Edge attributes are also embedded into 128 dimensions. The model consists of three Graph Transformer layers with 8 attention heads and a head dimension of 64. To enforce rotational equivariance, random global rotations are applied to molecular configurations during training as data augmentation (Abramson et al., 2024).
C.2 Training Configuration
For alanine dipeptide, models are trained using AdamW with weight decay $10^{-5}$ and a cosine learning-rate schedule ranging from $3\times 10^{-4}$ to $1\times 10^{-5}$, using a batch size of 256 for 5000 epochs. For MB potential, we use the same configuration except training for 2000 epochs. We find no consistent benefit from exponential moving average (EMA) of parameters and therefore do not use it in our experiments.
C.3 Inference
As Hutchinson’s trace estimator introduces bias for BGs (Tan et al., 2025a; Peng and Gao, 2025), we compute the divegence exactly using automatic differentiation. For inference, we use the Dormand–Prince 5(4) method (dopri5) with absolute and relative tolerances set to $10^{-4}$. Inference is performed with a batch size of 1000.
C.4 Computational Cost
The reported inference times (Tab. 4) correspond to generating $10^{4}$ samples. All other training and inference parameters are provided in §C.2 and §C.3. A batch size of 1000 is used for all experiments, except for the all-atom mapping, which uses a batch size of 512 due to GPU memory constraints.
Table 4: Training and inference time for CFM across settings. Inference times correspond to $10^{4}$ generated samples. Results marked with are from (Rehman et al., 2025).
Core Beta Heavy Atom All Atom ECNF++ SBG DiT CNF FALCON
Training 0.83h 1.39h 3.93h
Inference 0.84min 3.46min 13.92min
Total 0.84h 1.45h 4.16h 12.52h 16.83h 9.56h 7.65h
Appendix D Experimental Details for Force Matching
D.1 Architecture
Müller-Brown Potential. For the MB potential, we use a radial basis function (RBF) feature map followed by a multilayer perceptron. The RBF expansion has $K=100$ centers, initialized uniformly in $10,50]^{2}$ and optimized during training, with a fixed width $\sigma=5.0$. The features are passed through four fully connected layers of size 128 with softplus activation, followed by a linear output layer.
Alanine Dipeptide. For alanine dipeptide, the CG potential $U_{\eta}(\mathbf{R})$ is parameterized using the MACE architecture (Batatia et al., 2022), an equivariant message-passing graph neural network. Each CG bead is represented as a node in a geometric graph, with edges connecting neighbors within a cutoff radius and encoding relative position vectors. The model uses hidden irreducible representations of $32\times 0e+32\times 1o$, processed through two interaction layers with correlation order 3 and an angular momentum expansion truncated at $\ell_{\rm max}=3$. Node features are decoded by a readout layer ($16\times 0e$) into a scalar energy prediction ($1\times 0e$). Periodic displacement functions are applied during graph construction to handle boundary conditions correctly.
D.2 Training Configuration
For MB, training uses the Adam optimizer (via Optax) with a constant learning rate of $10^{-4}$ and batch size 128 for 500 epochs. For alanine dipeptide, training uses Adam with exponential learning rate decay ($\eta_{0}=10^{-3}$, decay rate 0.9), batch size 256, and 250 epochs. Training and validation splits are with a 90/10 ratio. Gradients are clipped to a global norm of 1.0. The final parameters for energy evaluation correspond to the minimum validation loss.
D.3 Computational Cost
The reported inference times (Tab. 5) correspond to evaluating $10^{4}$ samples. Training parameters follow §D.2. A batch size of 500 is used for inference.
Table 5: Training and inference time for MACE model. Inference times correspond to $10^{4}$ samples evaluated.
Core Beta Heavy Atom
Training 1.88h 4.11h
Inference 0.66s 0.72s
Appendix E Compute Infrastructure and Software
E.1 Hardware
All experiments, including model training, inference, and computational benchmarks, are performed on a single NVIDIA A100 GPU with 80 GB memory.
E.2 Software
CFM is implemented using JAX, Diffrax (Kidger, 2021), and Flax (Heek et al., 2024). Graph transformer is adapted from the implementation provided in https://github.com/noegroup/ScoreMD/blob/main/src/scoremd/models/graph_transformer.py. Training of CG PMF is carried out using chemtrain (Fuchs et al., 2025b) and chemtrain-deploy (Fuchs et al., 2025a) built on JAX, M.D. (Schoenholz and Cubuk, 2020). Molecular structures are visualized using OVITO (Stukowski, 2009).
Appendix F Algorithms
Algorithm 1 Training CG PMF via ESFM
Input: Dataset $\mathcal{D}_{\mathrm{bias}}=\{(\mathbf{r},\mathcal{F}_{\mathrm{proj}}(\mathbf{r}))\}$ (Rapidly converged); batch size $B$
Initialize: CG PMF network $U_{\eta}$
while not converged do
  Sample $\{(\mathbf{r}^{(i)},\mathcal{F}^{(i)}_{\mathrm{proj}})\}_{i=1}^{B}\sim\mathcal{D}_{\mathrm{bias}}$
  Compute $\mathbf{R}^{(i)}\leftarrow\Xi(\mathbf{r}^{(i)})$
  $\mathcal{L}_{\mathrm{ESFM}}\leftarrow\frac{1}{B}\sum_{i=1}^{B}\big\|\nabla_{\mathbf{R}}U_{\eta}(\mathbf{R}^{(i)})-\mathcal{F}^{(i)}_{\mathrm{proj}}\big\|_{2}^{2}$
  Update $\eta\leftarrow\mathrm{Optim}\big(\eta,\nabla_{\eta}\mathcal{L}_{\mathrm{ESFM}}\big)$
end while
Return $U_{\eta}$
Algorithm 2 Training Flow Model via CFM
Input: Dataset $\mathcal{D}=\{\mathbf{r}\}$ (biased or unbiased); batch size $B$
Initialize: flow model $v_{\theta}$
while not converged do
  Sample $\mathbf{r}^{(i)}\sim\mathcal{D}$
  Compute $\mathbf{R}_{1}^{(i)}\leftarrow\Xi(\mathbf{r}^{(i)})$
  Sample $\mathbf{R}_{0}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
  Sample $t^{(i)}\sim\mathcal{U}[0,1]$
  $\mathbf{R}_{t}^{(i)}\leftarrow(1-t^{(i)})\mathbf{R}_{0}^{(i)}+t^{(i)}\mathbf{R}_{1}^{(i)}$
  $u_{t}^{(i)}\leftarrow\mathbf{R}_{1}^{(i)}-\mathbf{R}_{0}^{(i)}$
  $\mathcal{L}_{\mathrm{CFM}}\leftarrow\frac{1}{B}\sum_{i=1}^{B}\big\|v_{\theta}(t^{(i)},\mathbf{R}_{t}^{(i)})-u_{t}^{(i)}\big\|_{2}^{2}$
  Update $\theta\leftarrow\mathrm{Optim}\big(\theta,\nabla_{\theta}\mathcal{L}_{\mathrm{CFM}}\big)$
end while
Return $v_{\theta}$
Algorithm 3 Sampling & Reweighting
Input: Trained models $U_{\eta}$, $v_{\theta}$; number of samples NN
 Initialize sample set $\mathcal{X}\leftarrow\emptyset$, log-weights $\mathcal{W}\leftarrow\emptyset$
for $i=1$ to NN do
  Sample $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
  Solve ODE $\frac{d\mathbf{R}}{dt}=v_{\theta}(t,\mathbf{R})$ with $\mathbf{R}(0)=\mathbf{z}$
  $\mathbf{R}^{(i)}\leftarrow\mathbf{R}(1)$
  $\Delta\ell^{(i)}\leftarrow\int_{0}^{1}\nabla\!\cdot v_{\theta}(t,\mathbf{R}(t))\,dt$
  $\log q_{\theta}(\mathbf{R}^{(i)})\leftarrow\log p_{0}(\mathbf{z})-\Delta\ell^{(i)}$
  $E^{(i)}\leftarrow U_{\eta}(\mathbf{R}^{(i)})$
  $\log\tilde{w}^{(i)}\leftarrow-\beta E^{(i)}-\log q_{\theta}(\mathbf{R}^{(i)})$
  Append $\mathbf{R}^{(i)}$ to $\mathcal{X}$ and $\log\tilde{w}^{(i)}$ to $\mathcal{W}$
end for
 Weight clipping for $\mathcal{W}$
 Normalize weights $\{w^{(i)}\}$ and compute ESS
Return samples $\mathcal{X}$, weights $\{w^{(i)}\}$, ESS
Appendix G Additional Results
G.1 Ramachandran Plots
We show Ramachandran plots for alanine dipeptide to illustrate the effect of importance reweighting. Fig. 6 shows the reference distributions from MD simulations in explicit and implicit solvent. Fig. 6b corresponds to a short 10 ns unbiased trajectory, which fails to sample the correct $\phi$ mode, highlighting the need for enhanced sampling to achieve sufficient configurational coverage.
We compare the MD reference with CG-BG flow proposals before and after reweighting for both mappings (Fig. 7, Fig. 8, Fig. 10, and Fig. 11). Unreweighted proposals exhibit noisy sampling and populate low-probability regions of the free energy landscape. These configurations receive low importance weights and contribute negligibly after reweighting, yielding distributions consistent with the reference.
We further examine reweighted samples obtained using different PMFs. The PMF trained on long unbiased data ($\text{PMF}_{U}$) underestimates the correct metastable basin of the $\phi$ angle (Fig. 9), reflecting the limitations of learning CG PMFs from slowly converging unbiased trajectories.
Figure 6: Ramachandran plots for MD simulations. (a) 500 ns unbiased MD simulation (explicit solvent). (b) 10 ns unbiased MD simulation (explicit solvent) . (c) 500 ns unbiased MD simulation (implicit solvent) (OBC2).
Figure 7: Ramachandran plots (Heavy Atom). (a) MD reference. (b) Flow proposal (flow trained on unbiased dataset). (c) Reweighted result ($\text{PMF}_{B}$).
Figure 8: Ramachandran plots (Heavy Atom). (a) MD reference. (b) Flow proposal (flow trained on WT-MetaD dataset). (c) Reweighted result ($\text{PMF}_{B}$).
Figure 9: Ramachandran plots (Heavy Atom). (a) MD reference. (b) Flow proposal (flow trained on unbiased dataset). (c) Reweighted result ($\text{PMF}_{U}$).
Figure 10: Ramachandran plots (Core Beta). (a) MD reference. (b) Flow proposal (flow trained on unbiased dataset). (c) Reweighted result ($\text{PMF}_{B}$).
Figure 11: Ramachandran plots (Core Beta). (a) MD reference. (b) Flow proposal (flow trained on WT-MetaD dataset). (c) Reweighted result ($\text{PMF}_{B}$).
G.2 Ablation of Weight Clipping
To stabilize importance reweighting, we apply a weight clipping strategy, discarding the top $1\%$ of samples with the largest log-weights.
As shown in Tab. 6, reweighting without clipping leads to large JS divergences and PMF errors, together with near-zero ESS, indicating severe weight degeneracy. In contrast, weight clipping restores stable estimates and substantially improves all metrics.
We further analyze sensitivity to the clipping ratio in Fig. 12. Aggressive clipping (10–20%) maximizes ESS but introduces bias in the reweighted distributions. Small clipping ratios (close to 0%) preserve unbiasedness but suffer from high weights variance. Balancing this bias–variance trade-off, we choose a conservative $1\%$ clipping threshold across all experiments, which yields stable metrics while maintaining physically consistent distributions.
Table 6: Quantitative comparison for alanine dipeptide across CG resolutions and different weight clipping thresholds, showing effects on JS, PMF errors, and ESS.
Model JS ($\downarrow$) PMF ($\downarrow$) ESS ($\uparrow$)
Flow Trained on Unbiased Dataset
Heavy Atom Proposal $0.0074\pm 0.0001$ $0.6051\pm 0.0102$
Heavy Atom Reweighted Proposal ($1\%$ clip) $0.0075\pm 0.0001$ $0.2493\pm 0.0126$ $0.2325\pm 0.0006$
Heavy Atom Reweighted Proposal ($0\%$ clip) $0.2364\pm 0.0584$ $8.3338\pm 3.2675$ $0.0000\pm 0.0001$
Core Beta Proposal 0.0072 $\pm$ 0.0001 $0.5236\pm 0.0099$
Core Beta Reweighted Proposal ($1\%$ clip) $0.0079\pm 0.0001$ $0.2519\pm 0.0136$ 0.2510 $\pm$ 0.0006
Core Beta Reweighted Proposal ($0\%$ clip) $0.2398\pm 0.0923$ $10.0688\pm 5.6594$ $0.0001\pm 0.0001$
Flow Trained on WT-MetaD Dataset
Heavy Atom Proposal $0.0505\pm 0.0002$ $2.5861\pm 0.0353$
Heavy Atom Reweighted Proposal ($1\%$ clip) $0.0090\pm 0.0002$ $0.2758\pm 0.0109$ $0.1936\pm 0.0005$
Heavy Atom Reweighted Proposal ($0\%$ clip) $0.1565\pm 0.0399$ $6.0221\pm 2.5237$ $0.0001\pm 0.0001$
Core Beta Proposal $0.0540\pm 0.0002$ $3.7156\pm 0.0393$
Core Beta Reweighted Proposal ($1\%$ clip) $0.0081\pm 0.0001$ 0.2325 $\pm$ 0.0093 $0.2239\pm 0.0005$
Core Beta Reweighted Proposal ($0\%$ clip) $0.5573\pm 0.1692$ $261.6973\pm 67.6421$ N/A
Figure 12: Metrics as a function of weight clipping ratio. (a) JS divergence. (b) PMF error. (c) ESS.
G.3 Free Energy of $\psi$ Dihedral
We provide additional free energy plots for the $\psi$ dihedral of alanine dipeptide (Fig. 13).
Figure 13: Comparison of $\psi$ dihedral free energy profiles across training settings. All panels show the MD reference alongside CG-BG proposals before and after reweighting. Results correspond to: (a) Flow trained on unbiased data (Core Beta); (b) Flow trained on WT-MetaD data (Core Beta); (c) Flow trained on unbiased (Heavy Atom); (d) Flow trained on WT-MetaD data (Heavy Atom).
G.4 Bond Length
Fig. 14 shows the C–N bond length distributions for the MD reference and CG-BG results before and after reweighting.
Figure 14: Comparison of C–N bond length distributions across training settings. All panels show the MD reference alongside CG-BG proposals before and after reweighting. Results correspond to: (a) Flow trained on unbiased data (Core Beta); (b) Flow trained on WT-MetaD data (Core Beta); (c) Flow trained on unbiased data (Heavy Atom); (d) Flow trained on WT-MetaD data (Heavy Atom).